(1-x 2 )(1-y)dx=xy(1+y)dy

(1-x2)(1-y)dx=xy(1+y)dy

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear upendra

(1-x2)(1-y)dx=xy(1+y)dy

seperate term

(1-x2)/x dx=y(1+y)/(1-y)  dy

(1/x  -x) dx={-(y+2)  +2/(1-y) }dy

now integrate

lnx -x2/2  =-(y+2)2/2  +-2ln|1-y|   +c


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin